OpenStudy (anonymous):

find the equation of the line tangent to the graph of h(x)=x^4-5x+2 at the point with x-coordinate 1.

OpenStudy (anonymous):

find the derivative; sub x=1 to get the slope; ;sub x=1 into h to get the point

OpenStudy (anonymous):

h'(x)=4x^3-5 slope is h'(1)=-1 y-coordinate is h(1)=...=-2, so point is (1,-2) y+2=-(x-1) or y=-x-1

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

:})